OpenStudy (anonymous):

What is the range of f(x) = log7x?

OpenStudy (anonymous):



OpenStudy (anonymous):

so is it all reals

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

domain is x>0, range is R

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Find the zero(s) of the function f(x) = log5x. Hint: The zeros are all the values of x when f(x) = 0.

OpenStudy (anonymous):

\[\log 5x=0\]\[10^{\log(5x)}=10^0\]\[5x=1\]\[x=\frac{1}{5}\]

OpenStudy (anonymous):

so the zeros are 1/5

OpenStudy (anonymous):

There is no value of log(x) that gives 0. No exponent to which any base can be raised that gives 0.

OpenStudy (anonymous):

@telliott99 incorrect. \[\log 5(\frac{1}{5})=\log 1=0\]because.\[1=10^0\]

OpenStudy (anonymous):

so the answer will be 0

OpenStudy (anonymous):

?

OpenStudy (anonymous):

do you understand what the "zero" of a function means?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so what is the zero of log (5x) ?

OpenStudy (anonymous):

1

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the zero of a function f(x) is the value x, such that f(x)=0

OpenStudy (anonymous):

log (5*1)=log 5 does not equal zero. so x=1 is not a "zero"

OpenStudy (anonymous):

so there is no solution

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@eseidl sorry my bad but No exponent to which any base can be raised that gives 0. so no x that log(x) = 0

OpenStudy (anonymous):

Again! I mean no p where e^p = 0

OpenStudy (anonymous):

@telliott99 your misunderstanding the nature of the logarithm. y=logx take "antilog" of both sides. \[10^y=x\]set y=0.\[10^0=x\]\[x=1\]what's the problem?

OpenStudy (anonymous):

I agree that there is no p such that \[e^p=0\]but this doesn't imply that there is no x such that\[\log_{a}x=0 \]If fact x=1 will give \[\log_{a}1=0 \]for any base a. Because,\[1=a^0\]

OpenStudy (anonymous):

Just tired and typing the wrong things. Apologies for causing confusion.